# Computer Vision Datasets and Models Exhibit Cultural and Linguistic Diversity in Perception

### v3

@@ Abstract
Computer vision often treats human perception as homogeneous: an implicit assumption that visual stimuli are perceived similarly by everyone. This assumption is reflected in the way researchers collect datasets and train vision models. By contrast, literature in cross-cultural psychology and linguistics has provided evidence that people from different cultural backgrounds observe vastly different concepts even when viewing the same visual stimuli. In this paper, we study how these differences manifest themselves in vision-language datasets and models, using language as a proxy for culture. By comparing textual descriptions generated across 7 languages for the same images, we find significant differences in the semantic content and linguistic expression. When datasets are multilingual as opposed to monolingual, descriptions have higher semantic coverage on average, where coverage is measured using scene graphs, model embeddings, and linguistic taxonomies. For example, multilingual descriptions have on average 29.9 more objects, 24.5 more relations, and 46.0 more attributes than a set of monolingual captions. When prompted to describe images in different languages, popular models (e.g. LLaVA) inherit this bias and describe different parts of the image. Moreover, finetuning models on captions from one language performs best on corresponding test data from that language, while finetuning on multilingual data performs consistently well across all test data compositions. Our work points towards the need to account for and embrace the diversity of human perception in the computer vision community.
@@ 1 Introduction
Computer vision researchers often make the implicit assumption that human perception is fundamentally homogeneous. In other words, they assume that—aside from supposedly small and, ultimately, unimportant variations—visual stimuli are perceived similarly by everyone. This homogeneity assumption is visible throughout the vision research field, from datasets to models to metrics. For instance, nearly no commonly used vision dataset, including CIFAR, ImageNet, MS-COCO, and PASCAL-VOC, contain information about their annotator’s background [34], suggesting that it is presumed that this background does not condition their visual perception [27, 68, 98].111Note that annotator information is included in datasets and models from other domains where the annotator’s background is thought to influence annotations, such as in NLP [87, 51] and medical vision [4, 35, 129]. Moreover, vision models trained on data from one demographic of annotators are taken to be representative of the perceptual modes of others across the world [76, 19]. Additionally, many common evaluation metrics average away the supposedly unimportant cross-annotator differences by measuring the mean model performance across multiple ground truth annotations [83, 6].
The homogeneity assumption stands against decades of research in cross-cultural psychology [36]. Early work in psychology suggests that fundamental aspects of visual perception, such as perception of length [101], geometrical intuition [84, 26], and depth [44] vary across people from different cultural and geographic backgrounds. Recent work has shown that differences in cultural backgrounds affect the way that individuals attend to, understand, and talk about visual content [80, 55, 15]. For instance, when looking at the same image, Americans tend to concentrate on the focal objects and its attributes, whereas Japanese tend to focus on the relationships between objects[74]. Other studies show that Westerners tend to perceptually group objects based on category membership, and Easterners based on resemblance [55].
Motivated by evidence from these psychological studies, we investigate if such differences in perception are detectable in vision datasets and model behaviors. We evaluate differences in perception across languages, using language as a rough proxy to culture — this proxy also being used in a strong history of studies on culture [33, 57]. According to the homogeneity assumption, the distribution of semantic content in the descriptions associated with images should be similar. If they are not similar, this would provide further evidence that perceptual biases are observable in vision datasets. Similarly, if models describe the same image differently when prompted to use different languages, we would have evidence that models have inherited such differences in their behavior.
By designing evaluations to examine image captions from datasets, vision-language models, and commercial API across languages, we show that perceptual differences are prevalent in vision datasets and affect model behaviors. In general, for a fixed set of images, multilingual annotations cover more semantic concepts and more expressive diversity than monolingual ones of the same size. Multilingual descriptions have on average more objects, more relations, and more attributes than a set of monolingual captions. Similarly, multilingual descriptions contain more expressive variety; for example, they contain more diversity in abstraction of concepts. These differences are reflected in model behavior: we find that models finetuned on captions translated from one language perform best when evaluated on its corresponding test set, indicating that models internalize the differences between captions in different languages. Overall, our primary contribution is a series of experiments which provide evidence against the homogeneity assumption by demonstrating the diversity (wide and substantive variation) of image annotations across languages and cultures in computer vision datasets and models.
Given these results, we propose several concrete recommendations for computer vision researchers. First, datasets should strive to include annotator information, such as cultural and geographical background. Secondly, multilingual vision models should be preferred over monolingual vision models, as merely translating captions does not guarantee representation of perceptual patterns in the target language. Thirdly, existing multilingual multimodal datasets [130, 98] may be exploited as rich sources of variation in human perception across the globe. Finally, designing computer vision systems explicitly for users from different viewpoints and cultures can improve both performance and accessibility.

### v4

# Semantic and Expressive Variation in Image Captions Across Languages

@@ Abstract
Most vision-language models today are primarily trained on English image-text pairs, with non-English pairs often filtered out. Evidence from cross-cultural psychology suggests that this approach will bias models against perceptual modes exhibited by people who speak other (non-English) languages. We investigate semantic and expressive variation in image captions across different languages; we analyze both human-annotated datasets and model-produced captions. By analyzing captions across seven languages (English, French, German, Russian, Chinese, Japanese, Korean) in high-quality image captioning datasets (Crossmodal and Visual Genome), we find that multilingual caption sets tend to provide richer visual descriptions than monolingual (including English-only) ones; multilingual sets contain more objects, more relationships, and more attributes. We observe the same results with multilingual captions produced by LLaVA and the Google Vertex API: for example, compared to monolingual captions, they cover more objects, more relations, and more attributes. These suggest that, across a large number of samples, different languages bias people and models to focus on different visual concepts. Finally, we show that models trained on image-text data in one language perform distinctly better on that language’s test set. Our work points towards the potential value of training vision models on multilingual data sources to widen the range/variation of descriptive information those models are exposed to.
@@ 1 Introduction
Typically, vision-language models are trained with large quantities of image-text pairs scraped from the web [116, 103, 33]. To obtain high quality image-text pairs, these web datasets are filtered with models like CLIP [94, 51] with the goal of ensuring that each text sample is sufficiently descriptive of or related to the corresponding image. However, this filtering tends to be biased towards English text samples, such that the majority of samples in popular image-text datasets are in English [37, 46, 103]. Many vision-language models are even trained only on English [94, 111, 71, 59], excluding other languages entirely.
Prior work shows that models tend to fail at tasks where naturally occurring text is uncommon, such as negation [30]. Therefore, increasing the lexical diversity of naturally occurring text is important: it allows models to experience a broader range of linguistic patterns that might be underrepresented within a certain set of scraped text, while maintaining the scale that is only possible through training on naturally occurring data. Thus, biasing vision datasets towards English-only image-text pairs may risk excluding the equally relevant and possibly unique visual concepts produced in languages other than English. Indeed, a wide body of work from the cross-cultural psychology suggests that people from across the world, speaking different languages and living in different cultural contexts, tend to describe the visual world differently. Early work in psychology suggests that fundamental aspects of visual perception, such as perception of length [106], geometrical intuition [89, 29], and depth [52] vary across people from different geographic backgrounds. More recent work has shown that individuals from different cultural backgrounds exhibit differences in how they look at, understand, and talk about visual scenes [85, 62, 16]. For instance, when looking at the same image, Americans tend to describe the focal objects and its attributes, whereas Japanese tend to focus on the relations between objects [80].
Drawing on this work, we hypothesize the content distribution of visual descriptions varies by the structure and use of specific languages – and that these differences are detectable in computer vision datasets and model behaviors. This hypothesis is further supported by observed differences in spoken language: German’s complex morphosyntactic system provides events with nuanced understanding of spatial relationships [92]; Russian verbs of motion require speakers to specify directionality [88]; English uses the generic verb of motion “to go” whereas a Russian speaker must specify whether the motion has a destination/purpose, the mode of transportation, and other possibly visually salient variables. Similar factors associated with different languages might encourage the expression of different visual concepts in descriptions. If true, our hypothesis suggests that models trained on English-only data would be further improved if exposed to a more global view.
In this paper, we investigate whether captions produced by people (in datasets) and by models tend to vary semantically and expressively across different languages. In our context, semantics refers to a caption’s content (what a caption “says”) as represented by scene graphs (objects, attributes, relations) [55, 63] and expression refers to how its content is communicated as measured by linguistic measures such as concreteness, authenticity, and tone [12]. We measure the variation of semantic and expression measurements between multilingual and monolingual sets of captions. If we systematically observe a greater variation in multilingual caption sets vs. monolingual sets (for example, if scene graphs created from multilingual caption sets are larger than monolingual ones), then this demonstrates that captions have different distributions of concepts in different languages. We analyze the human-produced captions from the high-quality Crossmodal dataset across seven diverse languages (English, French, German, Russian, Chinese, Japanese, Korean). The Crossmodal dataset keeps annotator instructions consistent across languages, ensuring fair comparison of multilingual and monolingual caption sets in our analysis by eliminating possible confounding sources of variation across languages.
Our results demonstrate that multilingual caption sets demonstrate higher semantic and expression variation than monolingual caption sets.
Importantly, while it is known that different people and models produce fairly widely varying captions in general [50, 9], our claim is that there are also broad distributional differences between captions produced in different languages, when evaluated over large numbers of samples.
Multilingual scene graphs are larger overall (that is, they cover more content) than monolingual scene graphs, with an increase of 46.0 objects, 66.1 relationships, and 66.8 attributes compared to scene graphs built from English-only captions.
Multilingual caption sets also exhibit broader expressive variation, such as an increase of 53.4 in range of tonality and 42.1 in coverage of the embedding space.
We extend this analysis to vision-language model behaviors (LLaVA and the Vertex API imagetext-001), and observe similar patterns in model-generated captions: multilingual scene graphs are larger than monolingual scene graphs overall by
21.9 objects, 18.8 relations, and 20.1 attributes, and multilingual caption sets have
63.0 wider range of tonality and nearly 92.4 wider coverage of the embedding space.
We further analyze how finetuning on linguistically diverse data affects a model’s captioning capabilities. Models are often finetuned on captions produced by humans (in datasets) or by other models (distillation). To understand this, we fine-tune models on captions from one language and evaluate them on captions from another language. Models finetuned on language perform significantly better on the test set from language , suggesting that models internalize language-specific distributional characteristics. For example, a model finetuned on captions translated from Japanese attains a SPICE F-score of .27 on reference captions translated from Japanese, but only .23 from English.
In sum, our primary contribution is providing evidence that captions produced by people and models vary by language. To do so, we give specific measurement methods on which multilingual distributions of captions have wider variation in content or information than monolingual distributions. Our work may help reframe the “curse of multilinguality” by emphasizing the diverse range of visual concepts in multilingual datasets. We are limited by our specific seven languages and mid-sized scale, but we believe our study provides a focused demonstration of the properties of multilingual data. Our work provides one possible explanation to findings in recent work [37] — that training on multilingual data at large scale improves vision representations — by understanding how the information in multilingual caption distributions might differ from monolingual ones, and relatedly, how monolingual captions differ across languages. See §6 for related work and §7 for limitations.
2 Measuring variation in datasets
We want to understand how captions in computer vision datasets vary across languages. Variation across languages in datasets may have downstream effects in associated benchmarks and downstream models. We choose to perform a close analysis of the Crossmodal (“XM”) dataset [115] because of its high quality and well-designed data annotation procedures. XM contains image descriptions in 36 languages over 3.6k geographically diverse images. To eliminate potential bias in captioning instructions across annotators (e.g., varying instructions in different languages), all annotators are provided with the same annotation instruction and required to be reading-proficient in English (in addition to being native/fluent in the target language). To ensure consistency in caption style, annotators across languages are primed with the same base caption. These procedures ensure that the XM dataset serves as a prime exemplar to fairly study differences between captions produced in different languages for the same sets of images.
Throughout the paper, we analyze captions produced in seven languages: English, French, German, Russian, Chinese, Japanese, and Korean. Together, these 7 languages encompass a wide range of typologically diverse linguistic families. Moreover, the speakers of these languages originate from a large variety of cultures and experiences.
2.1 Translation for fair comparison
We translate all captions into English with GPT-4 [86]
(gpt-4-0613)
to ensure fair comparison of caption content across languages by eliminating linguistic confounders.
Our measurements of semantic (§2.2) and expressive (§2.3) variation use tools like parsers, embedding models, and tokenizers which are either language-specific (e.g., entirely separate parsers for different languages [95, 17]) or language-biased (e.g., multilingual embeddings encode language-specific information [73, 87, 18, 91]).
When all captions are translated into English, we can analyze on the deeper differences in caption content and expression on “common ground”, without these linguistic confounders.
We chose GPT-4 because we can specify the manner and procedure for translation in the prompt to a degree not available for may other models; specifically, in our case it was more important for translations to be exact and preserve all factual details than to adhere to other translation values and metrics [15]. Several works have demonstrated GPT-4’s competence and control on multilingual tasks [99, 70, 54, 79, 129], particularly in capturing nuances of expression and meaning specific to a language.
We conduct a human evaluation to verify that the translations preserve salient conceptual information. We recruit 15 native speakers to rate 30 translations across each of the 6 non-English languages. Native speakers describe the translations as near perfect ( on a 1-5 scale, where 5 is “perfect” and 4 is “most information preserved”). Importantly, participants considered of important information in a visual scene (objects, spatial relations, color, etc.) as faithfully preserved across translation. See Appendix 8.1 for more details. We also verify statistically that variation among translated captions is not explained by the original language of the captions in Section 4. Participants were allowed to give open-ended feedback in the user study, and we find that even the most extreme corrections were minor – e.g. replacing “on” with “above” and “memories” with “impressions” – and therefore do not noticeably alter measurements. Therefore, we believe the translations did not introduce significant analysis-biasing artifacts.
Nonetheless, we repeat the representation variance experiment described in §2.3 and §3.4 with multilingual embeddings without translation and find that the same result holds as described (using English embeddings with translation) (see §9.2). This shows that our results still hold when using tools that can be theoretically applied across languages (even though they are language-biased). We describe all of our experiments in the main paper under the assumption of English translation for fair comparison.
2.2 Semantic variation across languages
We want to understand how captions across languages vary in their semantic content. A caption’s semantic content is what it says about the objects in the image and how they relate to each other. We ask: do captions across languages mention the same objects and relations, or are there systematic differences? Scene graphs are a common representation for this kind of information [55, 22, 53].
| Monolingual | Multilingual | ||||||||||||
| Metric | en | fr | de | ru | zh | ja | ko | avg | en-fr-zh | fr-zh-ru | de-fr-ru | avg | |
| Sem. | Objects | 2.59 | 2.92 | 3.16 | 3.03 | 2.99 | 3.41 | 2.71 | 2.98 | 3.71 | 3.92 | 3.93 | 4.35 |
| Relations | 1.54 | 1.76 | 1.94 | 1.88 | 1.71 | 1.99 | 1.59 | 1.77 | 2.41 | 2.57 | 2.57 | 2.94 | |
| Attributes | 1.27 | 1.66 | 1.97 | 1.74 | 2.01 | 2.47 | 1.46 | 1.78 | 2.36 | 2.59 | 2.55 | 2.97 | |
| Expr. | Concreteness | 1.80 | 2.24 | 2.11 | 2.03 | 2.26 | 2.25 | 2.14 | 2.12 | 2.08 | 2.13 | 2.10 | 2.17 |
| Authenticity | 33.21 | 40.09 | 42.22 | 32.35 | 34.02 | 33.83 | 35.28 | 35.86 | 53.12 | 54.41 | 53.35 | 53.21 | |
| Tone | 8.62 | 12.16 | 9.74 | 10.90 | 10.59 | 9.18 | 9.07 | 10.04 | 13.78 | 14.69 | 15.42 | 15.40 | |
| Embeddings | .38 | .42 | .43 | .40 | .46 | .42 | .43 | .42 | .54 | .54 | .49 | .52 | |
Method: Formally, a scene graph is defined as a list of tuples of the form or . Let , , and be the unique number of objects, relations, and attributes in a scene graph , respectively. These give measurements for the ‘size’ of . Let construct a scene graph from a caption , and represent the “union” of scene graphs — a new scene graph with shared content (objects, relations, attributes) only counted once. For each image out of all images in the dataset, there is a corresponding set of captions , where each caption may be denoted by , where is a language and is the caption index for captions in that language. Let be the set of languages. We create a monolingual set of captions for language : and a multilingual set of captions for a set of languages : . Let be a metric for scene graph size. We want to compare (mean size of scene graphs from monolingual captions) against (mean size of scene graphs from multilingual captions). If the latter is significantly larger than the latter, then we conclude that, there tend to be distributional non-overlaps/differences across the concepts covered in captions across languages (apart from differences across captions due to general annotator subjectivity). To implement which constructs a scene graph from a caption , we parse translated XM captions into scene graphs using FLAN-T5 [23] model fine-tuned on the FACTUAL-MR dataset. We find that LLM-based parsers were better able to resolve complex semantic relationships which often arose in the descriptions than conventional syntax-based parsers [105]. To implement the union of scene graphs, we concatenate their list representations and remove double counting of shared concepts (objects, relations, or attributes). We canonicalize the concepts using WordNet path similarity [81] and cosine similarity between concept embeddings [100] to merge concepts which have different text but refer to the same concept, erring on the side of merging. Multiple relationships between two entities are merged into a single edge.
Results: We find that multilingual unioned scene graphs cover more objects, relations, and attributes than monolingual ones (Table 1). For instance, unioning the scene graph from an English caption with the scene graphs from a French and a Chinese caption increase the resulting scene graph by objects, relations, and attributes. Further computing the sizes of intersections between monolingual unioned scene graphs shows that any two languages share only of objects and of relations in common on average (Table 11). Since image captions are short and targeted to focus on salient parts of the image, these observed differences between languages likely indicate different tendencies in what counts as salient across caption sets written by humans who speak different languages. Qualitatively, we observe that many of the objects represented in multilingual scene graphs but not in English scene graphs tend to be background objects and objects at varying levels of object composition (e.g. keyboard and CTRL key). Moreover, multilingual scene graphs tend to include more information about perspective (e.g. “zooming”, “fore/background”), color, size, and other details. See Figures 2 and 5 for detailed examples. Importantly, this demonstrates that there are distributional non-overlaps between the concepts represented in captions across languages.
2.3 Expressive variation across languages
We also want to understand how captions across languages vary by expression. Expression refers to how concepts are conveyed through language.
Method: We consider five linguistic measures of expression: Concreteness is computed across noun objects and indicates how much a word refers to a perceptible entity [13]. For instance, “purple” has a high concreteness rating (4.04 on a 5-point scale), whereas “beautiful” has a low one (2.16). Text is analytic when it expresses logical and hierarchical thinking patterns, demonstrates clout when it exhibits social status and confidence, is authentic when it is spontaneous, and varies by tone. These last four measures are provided by the Linguistic Inquiry and Word Counts (LIWC) framework [112], and are computed using word dictionaries and aggregation algorithms informed by linguistic and psychological research [12]. Many previous works have demonstrated the use of these measures in the quantification of emotional and psychological expression in text from a variety of contexts [102, 56, 1, 57]. A sixth measure is a model embedding, which captures some semantic but also important expressive dimensions of text [25, 20, 5].
To measure expressive variation, we compare the expressive coverage of multilingual caption sets and monolingual caption sets, where the expressive coverage of a set of text as measured by a measurement is , or the maximum width of spanned by elements of . Note that because embeddings are not one-dimensional, expressive coverage is measured as the maximum pairwise cosine distance between two points in a set. Like previously, we compare and . Following the interpretation methodology from §2.2, if the multilingual sets have more expressive coverage than monolingual sets, then this demonstrates some non-overlap in the distribution of captions’ expressive dimensions (as measured by the aforementioned metrics).
Results: Our results demonstrate variance in caption expression across languages. We find that across all metrics, expressive coverage is generally higher for multilingual rather than monolingual subsets (Table 1). For example, the range of tonality in caption sets widens by 53.4 for multilingual sets. We also find that the mean width of caption sets in model representations is larger for multilingual sets than monolingual sets. For example, the average coverage in embedding space of English XM captions increases from 0.38 to 0.54 when switching to an English, French, and Chinese caption set. This means that training with captions from multiple languages as opposed to one language may expose the model to a measurably wider range of caption expressiveness.
3 Measuring variation in model outputs
| Monolingual | Multilingual | |||||||||||||
| Model | Metric | en | fr | de | ru | zh | ja | ko | avg | en-fr-zh | fr-zh-ru | de-fr-ru | avg | |
| LLaVA | Sem. | Objects | 4.54 | 5.05 | 5.26 | 4.52 | 4.54 | - | - | 4.78 | 6.14 | 6.15 | 6.25 | 5.93 |
| Relations | 3.79 | 4.21 | 4.42 | 3.67 | 3.66 | - | - | 3.95 | 4.85 | 4.76 | 4.97 | 4.54 | ||
| Attributes | 2.75 | 3.47 | 3.50 | 2.76 | 3.25 | - | - | 3.15 | 4.00 | 3.99 | 4.07 | 3.86 | ||
| Expr. | Concreteness | 2.17 | 2.44 | 2.53 | 2.27 | 2.33 | - | - | 2.35 | 2.54 | 2.54 | 2.57 | 2.56 | |
| Authenticity | 35.97 | 38.01 | 34.33 | 37.20 | 44.63 | - | - | 38.03 | 54.94 | 54.64 | 54.82 | 53.15 | ||
| Tone | 6.79 | 9.53 | 16.64 | 16.48 | 11.56 | - | - | 12.20 | 16.74 | 20.79 | 19.80 | 16.56 | ||
| Embeddings | .22 | .29 | .28 | .29 | .36 | - | - | .29 | .45 | .47 | .43 | .47 | ||
| Vertex | Sem. | Objects | 3.65 | 3.51 | 3.60 | 3.86 | 3.46 | 3.13 | 3.18 | 3.48 | 4.13 | 4.24 | 4.13 | 4.17 |
| Relations | 2.96 | 2.83 | 2.89 | 3.20 | 2.68 | 2.37 | 2.47 | 2.77 | 3.45 | 3.48 | 3.38 | 3.40 | ||
| Attributes | 1.67 | 1.67 | 1.79 | 1.86 | 1.66 | 1.59 | 1.62 | 1.98 | 2.29 | 2.40 | 2.33 | 2.33 | ||
| Expr. | Concreteness | 1.64 | 1.64 | 1.67 | 1.66 | 1.51 | 1.50 | 1.56 | 1.60 | 1.75 | 1.74 | 1.73 | 1.81 | |
| Authenticity | 23.21 | 22.80 | 21.68 | 23.07 | 23.51 | 25.01 | 21.67 | 22.99 | 40.16 | 36.94 | 31.85 | 38.06 | ||
| Tone | 1.85 | 1.84 | 2.03 | 2.63 | 2.05 | 1.96 | 2.05 | 2.06 | 3.98 | 4.10 | 4.19 | 3.92 | ||
| Embeddings | .19 | .18 | .19 | .22 | .20 | .19 | .37 | .22 | .37 | .33 | .38 | .49 | ||
In §2, we demonstrated semantic and expressive variations in the XM dataset. These results make sense in light of social sciences studies that show how linguistic factors associated with different languages encourage the production of specific kinds of information. However, we also want to study if multilingual vision models also exhibit similar differences when producing captions in different languages. We ask the question: “what kinds of information differences might there be across captions produced by models in different languages for the same images?”
We study two models: LLaVA and the Google Vertex API. LLaVA [75] is a vision-language model capable of generating multilingual captions. Although LLaVA is trained with English data, it inherits multilingual capabilities from its large language model (LLM) component, LLaMA [119]. In total, we generate 54k captions for 3.6k Crossmodal images in 5 languages with 3 captions each, excluding Japanese and Korean for text quality issues. LLaVA allows us to study the significance of language factors across descriptions in an isolated fashion, because the vision representations are held constant across text prompts in different languages. See Appendix 8.2 for more details on probing multilingual behavior. The Google Vertex API is widely adopted to generate multilingual captions for images. The API outputs reflect established industry standards, having gone through extensive quality checks.
3.1 Semantic variation across languages
Method: We apply the same method as described in §2.2 to measure semantic variation, but substituting XM captions for LLaVA and Vertex-generated captions.
Results: We find that multilingual unioned scene graphs cover more objects, relations, and attributes than monolingual ones (Table 2). For instance, unioning the scene graph from an English caption with the scene graphs from a French and a Chinese caption increase the resulting scene graph by objects, relations, and attributes for LLaVA captions; and by objects, relations, and attributes for Vertex API captions. The smaller increase for Vertex API captions is likely explained by the more standardized (lower-variance) captions produced by Vertex API compared to LLaVA. Figure 2 displays an example of such a scene graph.
3.2 Can multilingual variation be explained by ‘different language modes’?
There may be a concern that multilingual captions generated in different languages by models are diverse because they are ‘(built to) operate in different modes’ for each language; for example, tokens from different languages might ‘activate’ language-specific representation axes [18], or APIs might route caption requests in different languages to model instances finetuned/trained on different data. Can all of the increase in concept coverage (as measured by larger scene graphs) for multilingual captions over monolingual captions be explained solely by models being built to work in ‘different modes’ for different languages? If the answer is ‘yes’, then the increase seems uninteresting. Although there are difficulties in measuring this, we make an argument that the answer is ‘no’. To do so, we compare multilingual captions produced by one model (LLaVA or Vertex API) against English captions produced by multiple models (e.g., LLaVA, GIT[121], BLIP [71]) and find that their variation/coverage to be similar. One would expect the variation among multilingual captions produced by one model to be substantially smaller than monolingual captions produced by different models, since the difference between caption variation due to ‘different language modes’ inside one model should be much smaller than the caption variation due to totally different training procedures, model architectures, etc. across different models that happen to be in the same language [27, 126]. However, if the variation between the two is roughly the same, then we cannot explain the variation among multilingual captions produced by one model to the model operating in ‘different language modes’.
Method: Formally, let be a caption produced by model for image in language . Let either be LLaVA or Vertex. For a general variation metric (e.g. size of unioned scene graphs or expressive coverage), we want to compare (variation metric across multilingual captions produced by the same model ) and (variation metric across English captions produced by different models). If these two values are similar, then the variation in LLaVA and Vertex’s multilingual outptus cannot be explained by models operating in ‘different language modes’.
Results: Our results suggest that demonstrated differences in scene graphs constructed from models’ captions are not entirely explained by ‘different language modes’. Interestingly, we find that the same-model across-language scene graphs are only slightly smaller than the same-language across-model scene graphs: objects, relations, and attributes. See Table 12 for more details. This supports the significance of caption content differences across languages. Furthermore, we find that caption sets generated by one model (Vertex) in different languages have large proportions of concepts not covered by English GIT and BLIP models: of objects, of relations, and of attributes. See Table 13 for more details.
| Monolingual | Multilingual | ||||||||||||
| en | de | fr | ru | zh | ja | ko | en,fr,zh | fr,ru,zh | de,fr,ru | avg | |||
| Vertex | 9.6% | 9.7% | 9.9% | 10.3% | 9.8% | 9.9% | 9.3% | 11.7% | 11.7% | 11.2% | 11.5% | ||
| 1.55 | 1.60 | 1.62 | 1.69 | 1.60 | 1.59 | 1.49 | 1.92 | 1.92 | 1.84 | 1.90 | |||
| LLaVA | 9.5% | 10.9% | 11.6% | 9.3% | 9.1% | - | - | 12.0% | 12.1% | 12.5% | 12.0% | ||
| 2.86 | 3.30 | 3.55 | 2.83 | 2.75 | - | - | 3.71 | 3.70 | 3.86 | 3.69 | |||
3.3 Grounding variation in scene graphs
Another concern may be that captions generated by a model might contain concepts that are not ‘objectively’ grounded in the image. For example, one caption might mention a ‘tree’, another might mention ‘tree trunk’ and ‘branches’, possibly to describe specific attributes of each (e.g., the ‘tree trunk is brown, and the branches are black’) or describe a unique relation between them (e.g., ‘the branches are all on the same side of the tree trunk’). Although these types of differences are important, since they reflect complementary information about the content of the image, if variation across caption content and expression can be explained solely by variation induced by these kinds of ‘subjective’ focuses, then one could argue it is not as significant or interesting. Therefore, we reproduce our scene graph experiment to measure content variation only on object sets documented in the Visual Genome object list [63] for that image. Since the Visual Genome object list is standard and shared across all models, it provides an ‘objective’ grounding for measuring the size of scene graphs by objects.
Method: Formally, let be the Visual Genome object set for an image . Recall from §3.1 that and are the monolingual and multilingual caption sets using language and language set , respectively, for image . Let take in a caption set and output the unique objects (as if constructing a unioned scene graph and then extracting only the objects). We want to compare ( unique VG objects covered by monolingual caption set) and ( unique VG objects covered by multilingual caption set). If the latter is significantly larger than the former, then this is a stronger confirmation that there tend to be distributional non-overlaps/differences across concepts covered in captions across languages.
Results: The results demonstrate, like before, that multilingual scene graphs capture more Visual Genome objects/concepts than monolingual scene graphs (see Table 3). For example, unioning an English scene graph with a French and Chinese scene graph captures more Visual Genome-annotated objects than unioning with two other English scene graphs. Interestingly, we also find some evidence that multilingual scene graphs are able to identify objects which are unidentified even in Visual Genome’s dense annotations (App. Table 16).
3.4 Expressive variation across languages
Method: We apply the same method as described in §2.3 to measure semantic variation, but substituting XM captions for LLaVA and Vertex-generated captions.
Results: Our results demonstrate variance in models’ caption expression across languages. We find that across all metrics, expressive coverage is generally higher for multilingual rather than monolingual subsets (Table 2). For example, the range of noun concreteness in caption sets widens by for multilingual sets on average for LLaVA, and by for Vertex. We also find that the mean width of caption sets in model representations is larger for multilingual sets than monolingual sets. For example, the average coverage in embedding space of English LLaVA captions increases from 0.22 to 0.45 when switching to an English, French, and Chinese caption set, and from 0.19 to 0.37 for Vertex.
4 Effects of Fine-tuning Across Language
In §2 and §3, we showed that dataset and model captions exhibit semantic and expressive variation across languages. What effect does this have on how we train and fine-tune captioning models? Do models internalize the specific distributional characteristics of captions from one language?
Method: To evaluate this, we finetune a pretrained vision-language model, GIT [121] – chosen for its simple architecture and robust performance across benchmarks, on 8 constructed image caption sets. These finetuning datasets use the same set of 1.8k training images from the Crossmodal dataset (“XM”) [114]. The first 7 training sets contain translations from a single language (e.g. 1.8k captions all from French, 1.8k captions all from Chinese). The eighth dataset is constructed with equal proportions of captions across all 7 languages. We inference the finetuned models on a hold-out set of 1.8k images from XM and evaluate performance against each of the the corresponding 8 sets of image caption labels. We use SPICE F1-score [4], which measures conceptual overlap over low-level syntax adherence, as our evaluation metric. Validating each of the 8 fine-tuned models on each of the 8 holdout caption sets yields 64 evaluations. We perform this entire process on captions from XM, LLaVA, and Vertex.
We interpret our results using the following logic. Let represent the score of a model fine-tuned on the data from language on the evaluation data from language (suppose bigger is better). If we observe superior model test performance on the test set corresponding to the language it was trained on relative to other test sets (that is, ), the model has internalized distributional characteristics of each language.
Results: We find that a model fine-tuned on language performs best on language (Table 4) for Vertex captions. We find the same results across models finetuned on data from LLaVA (Table 14), and from the human annotated XM dataset (Table 15), including when evaluating with default loss rather than SPICE F-score. Moreover, finetuning on a multilingual collection of captions yields a model which performs consistently well across all evaluation data compositions. In the case of training on human-generated multilingual captions (from the Crossmodal dataset), this model is the second-best model for 5/7 holdout-set languages out of 8 models, coming only behind a model finetuned on that language itself. Together, these suggest that models trained with multilingual data may represent diverse perceptual modes and learn more about the visual world.
| Evaluated on | |||||||||
| en | de | fr | ru | zh | ja | ko | multi | ||
| en | 0.225 | 0.229 | 0.219 | 0.218 | 0.229 | 0.232 | 0.230 | ||
| de | 0.213 | 0.219 | 0.217 | 0.215 | 0.210 | 0.226 | 0.219 | ||
| fr | 0.248 | 0.240 | 0.234 | 0.236 | 0.239 | 0.253 | 0.246 | ||
| ru | 0.226 | 0.234 | 0.228 | 0.231 | 0.236 | 0.237 | 0.239 | ||
| zh | 0.199 | 0.202 | 0.199 | 0.207 | 0.220 | 0.224 | 0.216 | ||
Fine-tuned on |
ja | 0.212 | 0.212 | 0.215 | 0.212 | 0.226 | 0.245 | 0.223 | |
| ko | 0.218 | 0.222 | 0.224 | 0.217 | 0.242 | 0.239 | 0.235 | ||
| multi | |||||||||
5 Discussion
5.1 Implications
Reframing the “curse of multilinguality”. We contribute another framing to the ongoing debate related to the “curse of multilinguality.” On one hand, arguments in favor of multilingual language and vision models are often made from an accessibility perspective of expanding information and technology access to individuals [47]. Meanwhile, empirical evidence suggests that increasing the number of languages a model handles degrades monolingual performance [26]. Multilingualism is, seen as a trade-off against model performance. Our work provides support for the possibility of a system benefit from including multilingual data. Using “organically” / human-produced multilingual data introduces substantive changes in the semantic concepts mentioned for the same images. Thus, incorporating the diverse content distribution may improve model robustness [58, 96, 35, 37]. Translating English content to other languages may be missing some information. Many efforts to build vision-language models for a non-English language train on translations from English to [82, 19, 41, 36, 90, 93]. Although these systems are usable in , our investigation suggests that they may miss out on the unique characteristics of an “organically ” content distribution (i.e., that native -speakers would provide). Indeed, other works which collect captions from -speakers and train models on them suggest quantitative and qualitative improvements over training from English-to- translated captions [38, 125]. Our work provides an explanation for such observations. Wariness of a “perceptual monoculture”. The English-dominated language modeling landscape may not only introduce accessibility issues for non-English speakers but also possibly a “perceptual monoculture” [60]. In these ways, our contribution shares motivation and aims with a diverse array of previous works in how models and datasets encode and internalize particular limiting and possibly harmful knowledge, such as by gender [14], nationality [101], and language [122, 2].
5.2 Recommendations
Multilingual modeling from organically multilingual data. Currently, many multilingual vision-language models are built by training models on captions translated from English into the desired language. When possible, multilingual vision-language applications should be built by training models collected from native speakers. Our results suggest that the advantage is not only syntactic (i.e. more “natural-sounding” than translations), but that annotations produced in the desired language have specific distributional semantic and expressive characteristics.
Multilinguality as visual descriptive diversity. Multilingual vision-language datasets should be explored as repositories of perceptually diverse annotations. At a basic level, simply training models on multilingual rather than monolingual vision datasets may expose them to a more diverse, rich set of concepts. At a more complex level, specifically designed structured learning paradigms such as contrastive learning may be able to exploit concept diversity across multilingual captions to yield richer image representations.
5.3 Conclusion
This paper demonstrates semantic and expressive in image captions across languages, revealing meaningful differences in how visual content is represented. Our results support that training models on multilingual captions exposes them to a wider range of semantic and expressive material. This increased range may result in more robust vision representations [37]. By embracing the richness that different languages bring, we can move toward more inclusive, nuanced models that better reflect the diversity of human perception.
Acknowledgements
We thank Katharina Reinecke for the many insightful discussions that inspired this work; Quan Ze (Jim) Chen, Matthew Wallingford, Sarah Pratt, Thao Nguyen, Hyunwoo Kim, Wei-Chiu Ma, Chris Lin, Lisa Orii, Nino Migineishvili, Ashish Sharma, Jae Sung (James) Park and members of RAIVN lab for the multiple rounds of discussion, feedback and reviewing; Cheng-Yu Hsieh and Ashish Sharma for code/package assistance; Allen AI for running the surveys; and Chun-Liang Li at Google Cloud for running the Vertex API. We also thank participants of our study for carefully captioning the images.
Supplementary Material